# Causally constrained reduced-order neural models of complex turbulent dynamical systems

### v2

@@ Acknowledgements.
I Fluctuation-Dissipation Theorem (FDT) from data: quasi-Gaussian and score matching
@@
Consider a $n$-dimensional system $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$. The Fluctuation-Dissipation Theorem (FDT) states that the time-dependent response of an observable $A(x_{k}(t))$ to a small impulse perturbation $\delta x_{j}(0)$ imposed on $x_{j}(0)$ at time $t=0$, can be retrieved solely from stationary statistics of the system as:
$R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}=-\Big\langle x_{k}(t)\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}\Big\rangle.$ (1)
where $\rho(\mathbf{x})$ represents the invariant probability distribution of the system [30, 74]. The FDT establishes a link between perturbed quantities and stationary statistics, i.e. the invariant measure of the system [63, 74]. The brackets $\langle\cdot\rangle$ represents ensemble averages, in practice computed using temporal averages under the assumption of ergodicity. Application of the FDT in the main text focused on the observable $A(x_{k}(t))=x_{k}(t)$. In this case, Eq. 1 represents the impulse response in the ensemble mean.
I.1 The FDT from data
@@
The score-matching procedure introduced by Hyvärinen (2005) [45] provides an elegant way to estimate the score $s(\mathbf{x})$ directly on the system’s inertial manifold. In practice, however, this estimation can be prone to several sources of error. First, ensemble averages must be replaced by empirical averages computed from finite datasets. Second, the inferred score inevitably depends on the chosen neural-network architecture and its associated hyperparameters. Nevertheless, the theoretical formalism of the FDT, provides analytical constraints on the instantaneous response $\mathbf{R}(t=0)$. As shown in [37], these constraints can be used as a posteriori correction of the estimated score, thereby ensuring reliable results. Such corrections are non-unique, but they typically produce only small adjustments to the inferred response and thus eliminate the need for more sophisticated architectures. In what follows, we detail the analytical correction for score-based estimates relevant to the ensemble-mean response. We denote the ideal score by $s(\mathbf{x})$, the score inferred through score matching by $s_{\epsilon}(\mathbf{x})$, and the corrected score by $s_{c}(\mathbf{x})$. The system’s dynamics is encoded in a data matrix $\mathbf{x}\in\mathbb{R}^{T,n}$, where $T$ is the number of temporal samples and $n$ the system’s dimensionality. The correction procedure is dependent on the type of response analyzed. If the interest is in the response in ensemble mean, i.e. observable $A(x_{k}(t))=x_{k}(t)$ in Eq. (1), then it is possible to use the correction already outlined in [37]. The response operator in the ensemble mean at time $t=0$ of any system is:
$\mathbf{R}(0)=\mathbf{I}$ (4)
$\mathbf{I}$ being the identity matrix. The empirical estimation of the instantaneous response operator with the inferred score $s_{\epsilon}(\mathbf{x})$ will lead to small errors, so that
$\mathbf{R}_{\epsilon}(0)=-\frac{1}{T}\mathbf{x}^{\text{T}}\cdot s_{\epsilon}(\mathbf{x})=\mathbf{I}-\boldsymbol{\epsilon}$ (5)
where $\mathbf{R}_{\epsilon}(0)$ denotes the estimated response. It is then possible to correct the inferred score as:
$s_{c}(\mathbf{x})=s_{\epsilon}(\mathbf{x})\cdot(\mathbf{I}-\boldsymbol{\epsilon})^{-1}.$ (6)
@@
Figure 2: Time-dependent, responses in ensemble mean $x_{k}(t)$ to an impulse perturbation in the CdV model (Eq. (3) in the main text) as predicted by the FDT using the quasi-Gaussian and score-matching approximations. Example: column (1), row (2) quantifies the time-dependent mean response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$.
II Possible limitations of the causal constraints and practical considerations
Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [29]. Specifically, the response operator at the shortest time-scale can be written in terms of the ensemble averaged Jacobian as $\mathbf{R}(dt)=\langle\mathbf{J}(\mathbf{x})\rangle dt+\mathbf{I}$, where $\mathbf{J}(\mathbf{x})$ denotes the Jacobian, $\mathbf{I}$ is the identity matrix, and $dt\rightarrow 0$. Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss.
A complication can in principle arise in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [57, 58]. In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise. One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches advocated in [22]. For example, Ref. [76] proposed a nonlinear coordinate transformation to quotient the $Z_{2}$ symmetry of the Lorenz ’63 system [57], leading to the so-called “proto-Lorenz” system. In the case of the Lorenz ’96 system, defined as $\dot{u_{j}}=(u_{j+1}-u_{j-2})u_{j-1}-u_{j}+F,~j=0,...,J$ with periodic boundary conditions, the model is translationally invariant, with each $u_{j}$ having the same ensemble mean $\langle u_{j}\rangle=\bar{u}$. In this latter case, introducing a site-dependent forcing $F_{j}$ is sufficient to guarantee the validity of our framework. For example, one could define $F_{j}=F+\eta_{j}$, with $\eta_{j}\sim\mathcal{N}(0,\sigma^{2})$, mimicking the effect of topography, and therefore leading to a more realistic forcing term.
Crucially, we argue that the limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems, with no symmetries. As discussed in [29], in the case of climate modeling, the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations. In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse-graining. From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
$d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$
in which the structural constraint $\partial f_{k}(\mathbf{x})/\partial x_{j}=0$ is explicitly enforced. The validity of such constraints is then assessed by the accuracy of the resulting stationary and forced responses, rather than by comparison to a non-existent unique underlying dynamics.
@@ II Possible limitations of the causal constraints and practical considerations
Finally, causally constrained emulators can be readily compared with an unconstrained counterpart by evaluating their representation of the linear response operator against the FDT-based estimate in Eq. (1) of the main text. Such comparisons should be performed at short time lags to avoid spurious responses that may arise at longer lags when using the FDT [28]. The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model. In systems with exact symmetries, as discussed above, any biases introduced by the constraints will be immediately apparent in the representation of stationary variability.
III Details on the emulator formulations
We report the main technical details for fitting the emulators in the main text. This Section follows Appendix C in [29] and we refer to that work for more details. Consider a $n$-dimensional dynamical system described by a stochastic differential equation $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$. This equation can be discretized by the forward Euler method as follows:
[…]
where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$. Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [54, 53, 51, 99]. The strategy is then to fit a rescaled version of Eq. (7) from observational data alone. Such rescaled equation takes the following form:
$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t).\end{split}$ (8)
In what follows, we will not distinguish between the original model terms (e.g., $\mathbf{f}$) and their emulator counterparts, in order to avoid unnecessary notation and maintain a clear and concise presentation. We explicitly clarify the distinction whenever it becomes relevant. We proceed in two steps: fitting the determistic drift and then the stochastic noise. We first focus on fitting the drift. We compute $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)$ and fit the deterministic dynamics such that $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))$. In the main paper, we considered two cases:
  1. i.
    Unconstrained drift. The vanilla emulator is trained to minimize the MSE between $\Delta\mathbf{x}$ and $\mathbf{f}(\mathbf{x}(t))$ as:
    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))~.$ (9)
    The model $\mathbf{f}$ is fitted by a multilayer perceptrong (MLP). The MSE is minimized using a version of the gradient descent algorithm (Adam, [48]) with weight decay. The MLP considered here always consisted of one hidden layer with 100 neurons. We found empirically that adding more layers can lead to instabilities even in stationary runs, while opting for a shallow, but wide network led to stable simulations. The activation function used is the swish function (SiLU), which we found yields smoother responses to external perturbations.
  2. ii.
    Causally constrained drift. Here we consider the same architecture of the unconstrained drift above. The main difference is in the loss function, here equal to:
    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (10)
    where the set of entries $(k,j)$ is determined by the causal Adjacency matrix introduced in the main text. The derivative $\partial f_{k}/\partial x_{j}$ are computed with Autograd from PyTorch, providing automatic differentiation [80]. $\lambda$ is here set to $\lambda=10$.
Once a determistic drift $\mathbf{f}$ is computed we close the equation with a Wiener process. The first step is to compute the residuals of the deterministic dynamics as $\mathbf{r}(t)=\Delta\mathbf{x}-\mathbf{f}(\mathbf{x}(t))$, $\mathbf{r}(t)\in\mathbb{R}^{n}$. A stochastic parametrization for this residual term is then constructed. As the main paper is focused on proposing constraints acting on the drift term, we adopted the simplest option, in which the noise term in Eq. (8) is diagonal, with each entry $(i,i)$ given by the standard deviation of $r_{i}(t)$. For details on more complex stochastic closure, such as state-dependent noise we refer to the work of Falasca (2025) [29].
IV Charney De-Vore model: parameters
@@
V Practical estimation of the response operator
@@
Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$, we build the impulse response operator as follows:
@@
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $\Delta_{j}$ to the degree of freedom $x_{j}$ at time $t=0$. The amplitude of the perturbation should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems. In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
@@
    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$. This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
@@
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$. The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
@@
    We define the impulse response operator for observable $A(x_{k}(t))$ as
    $R_{k,j}(t)=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
VI Neural emulators of the Charney-DeVore model: stationary statistics
@@
Figure 3: Stationary statistics. First and second rows: stationary distributions (PDFs) of the variables $x_{1},...,x_{6}$ of the CdV model. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator. Third and fourth rows: same as the first row but for the autocorrelation functions (ACFs).
VII Response operator for the Charney-DeVore model
@@
Figure 4: Time-dependent, responses in ensemble mean of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
@@
Figure 5: Time-dependent, responses in ensemble variance of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
@@
Figure 6: Perturbed statistics in the linear regime. Panel (a): Total MSE over time for the mean response, computed from the full response operator $\mathbf{R}(t)\in\mathbb{R}^{6,6}$ for each model. Panel (b): Same as Panel (a) but the response in ensemble variance.
@@
VIII Response to large step function forcings in the Charney-DeVore model

### v3

@@ II.2 Data-driven emulation of the Charney-DeVore model from subsampled data
III Fluctuation-Dissipation Theorem (FDT) from data: quasi-Gaussian and score matching
@@
Consider a $n$-dimensional system $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$. The Fluctuation-Dissipation Theorem (FDT) states that the time-dependent response of an observable $A(x^{(k)}_{t})$ to a small impulse perturbation $\delta x^{(j)}_{0}$ imposed on $x^{(j)}_{t=0}$ at time $t=0$, can be retrieved solely from stationary statistics of the system as:
$R^{k,j}_{t}=\lim_{\delta x^{(j)}_{0}\to 0}\frac{\delta\langle A(x^{(k)}_{t})\rangle}{\delta x^{(j)}_{0}}=-\Big\langle A(x^{(k)}_{t})\frac{\partial\ln\rho(\mathbf{x})}{\partial x^{(j)}}\Big|_{\mathbf{x}_{0}}\Big\rangle.$ (24)
where $\rho(\mathbf{x})$ represents the invariant probability distribution of the system [24, 66]. The FDT establishes a link between perturbed quantities and stationary statistics, i.e. the invariant measure of the system [56, 66]. The brackets $\langle\cdot\rangle$ represents ensemble averages, in practice computed using temporal averages under the assumption of ergodicity. Application of the FDT in the main text focused on the observable $A(x^{(k)}_{t})=x^{(k)}_{t}$. In this case, Eq. 2 represents the impulse response in the ensemble mean.
III.1 The FDT from data
@@
The score-matching procedure introduced by Hyvärinen (2005) [39] provides an elegant way to estimate the score $s(\mathbf{x})$ directly on the system’s inertial manifold. In practice, however, this estimation can be prone to several sources of error. First, ensemble averages must be replaced by empirical averages computed from finite datasets. Second, the inferred score inevitably depends on the chosen neural-network architecture and its associated hyperparameters. Nevertheless, the theoretical formalism of the FDT, provides analytical constraints on the instantaneous response $\mathbf{R}_{t=0}$. As shown in [31], these constraints can be used as a posteriori correction of the estimated score, thereby ensuring reliable results. Such corrections are non-unique, but they typically produce only small adjustments to the inferred response and thus eliminate the need for more sophisticated architectures. In what follows, we detail the analytical correction for score-based estimates relevant to the ensemble-mean response. We denote the ideal score by $s(\mathbf{x})$, the score inferred through score matching by $s_{\epsilon}(\mathbf{x})$, and the corrected score by $s_{c}(\mathbf{x})$. The system’s dynamics is encoded in a data matrix $\mathbf{x}\in\mathbb{R}^{T,n}$, where $T$ is the number of temporal samples and $n$ the system’s dimensionality. The correction procedure is dependent on the type of response analyzed. If the interest is in the response in ensemble mean, i.e. observable $A(x^{(k)}_{t})=x^{(k)}_{t}$ in Eq. (24), then it is possible to use the correction already outlined in [31]. The response operator in the ensemble mean at time $t=0$ of any system is:
$\mathbf{R}_{0}=\mathbf{I}$ (27)
$\mathbf{I}$ being the identity matrix. The empirical estimation of the instantaneous response operator with the inferred score $\tilde{s}(\mathbf{x})$ will lead to small errors, so that
$\tilde{\mathbf{R}}_{0}=-\frac{1}{T}\mathbf{x}^{\text{T}}\cdot\tilde{s}(\mathbf{x})=\mathbf{I}-\boldsymbol{\epsilon}$ (28)
where $\tilde{\mathbf{R}}_{0}$ denotes the estimated response. It is then possible to correct the inferred score as:
$s_{c}(\mathbf{x})=\tilde{s}(\mathbf{x})\cdot(\mathbf{I}-\boldsymbol{\epsilon})^{-1}.$ (29)
@@
Crucially, depending on the data, it may be useful to compute the response $R^{k,j}_{t}$ on standardized data. This is especially true when variables span disparate scales, as in principal component representations.
@@
Figure 5: Time-dependent, responses in ensemble mean $x^{(k)}_{t}$ to an impulse perturbation in the CdV model (Eq. (3) in the main text) as predicted by the FDT using the quasi-Gaussian and score-matching approximations. Example: column (1), row (2) quantifies the time-dependent mean response of $x^{k=2}_{t}$ given a small, impulse perturbation imposed on $x^{j=1}_{0}$ at time $t=0$.
IV Possible limitations of the causal constraints and practical considerations
Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [23]. Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x^{(j)}$ in the neural network loss.
A complication can in principle arise in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [50, 51]. In such cases $R^{k,j}_{1}=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x^{(j)}\neq 0$ pointwise. One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches advocated in [18]. For example, Ref. [68] proposed a nonlinear coordinate transformation to quotient the $Z_{2}$ symmetry of the Lorenz ’63 system [50], leading to the so-called “proto-Lorenz” system. The limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems. In these realistic flows, symmetries are always broken.
@@ IV Possible limitations of the causal constraints and practical considerations
Crucially, as also noted in the main text, differently from the rigid (data independent) physical constraints, the utility of causal constraints depends on the system under study and on the accuracy of the data-driven causal discovery. The benefit of augmenting physics-constrained models with causal information can be assessed directly from data by comparing the FDT-based operator with the emulators’ response operators. Such comparisons should be performed at short time lags to avoid spurious responses that may arise at longer lags when using the FDT [22]. The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model. In systems with exact symmetries, as discussed above, any biases introduced by the constraints will be immediately apparent in the representation of stationary variability.
V Charney De-Vore model: parameters
@@
VI Practical estimation of the response operator
@@
Given a model $\mathbf{x}_{t+1}=\mathbf{f}(\mathbf{x}_{t})+\mathbf{\Sigma}\bm{\xi}(t)$, we build the impulse response operator as follows:
@@
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $\Delta_{j}$ to the degree of freedom $x^{(j)}$ at time $t=0$. The amplitude of the perturbation should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems. In practice we do as follows: we consider the long time series of $x^{(j)}_{t}$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x^{(j)}_{t}$.
@@
    Therefore, for a given initial condition $\mathbf{x}_{0}$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$. This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
@@
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x^{(k)}_{t})\rangle$ of an observable $A(x^{(k)}_{t})$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x^{(k)}_{t})\rangle_{\text{p}}$ and $\langle A(x^{(k)}_{t})\rangle$. The observables considered in this study are going to be: (i) $A(x^{(k)}_{t})=x^{(k)}_{t}$ and (ii) $A(x^{(k)}_{t})=(x^{(k)}_{t}-\mu^{(k)}_{t})^{2}$, with $\bm{\mu}_{t}$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
@@
    We define the impulse response operator for observable $A(x^{(k)}_{t})$ as
    $R^{k,j}_{t}=\frac{\langle A(x^{(k)}_{t})\rangle_{\text{p}}-\langle A(x^{(k)}_{t})\rangle}{\Delta_{j}}.$
VII Neural emulators of the Charney-DeVore model: stationary statistics
@@
Figure 6: Stationary statistics. First and second rows: stationary distributions (PDFs) of the variables $x^{(1)},...,x^{(6)}$ of the CdV model. Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information. Third and fourth rows: same as the first row but for the autocorrelation functions (ACFs).
VIII Response operator for the Charney-DeVore model
@@
Figure 7: Time-dependent, responses in ensemble mean of variable $x^{(k)}_{t}$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x^{(k=2)}_{t}$ given a small, impulse perturbation imposed on $x^{(j=1)}_{0}$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
@@
Figure 8: Time-dependent, responses in ensemble variance of variable $x^{(k)}_{t}$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x^{(k=2)}_{t}$ given a small, impulse perturbation imposed on $x^{(j=1)}_{0}$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
@@
Figure 9: Perturbed statistics in the linear regime. Panel (a): Total MSE over time for the mean response, computed from the full response operator $\mathbf{R}_{t}\in\mathbb{R}^{6,6}$ for each model. Panel (b): Same as Panel (a) but the response in ensemble variance.
@@
IX Response to large step function forcings in the Charney-DeVore model